# LaViDa: A Large Diffusion Language Model for Multimodal Understanding

### v1

@@ A.1 Formulation of Discrete Diffusion Models (DMs)
In this section, we provide a detailed review of the formulation of Discrete Diffusion Models (DMs) briefly described in 2.2. Given a text sequence of $L$ tokens $X_{0}=[X_{0}^{1},X_{0}^{2},...X_{0}^{L}$, the forward process $q(X_{t}|X_{s})$ gradually coverts it to a sequence of mask tokens $[M]$, denoted by $X_{1}=[X_{1}^{1},X_{1}^{2},...X_{1}^{L}]$ over the continuous time-interval $[0,1]$, with $1\geq t\geq s\geq 0$. Formally, this process is defined as
@@ A.2 Timestep shifting Schedule
Rounding in Discretization. In principle, we can pick any $f(.)$. However, given a particular choice of $L$ and $K$, if $\lfloor f(t_{k})L\rfloor$ and $\lfloor f(t_{k-1})L\rfloor$ yields the same integer, then no tokens are unmasked when we compute $p_{\theta}(X_{t_{k-1}}|X_{t_{k}})$. Hence, the actual number of function calls to the model $\theta$ may be less than, $K$ depending on the choice of $f(.)$. To address make the sampling compute cost more predictable and allow for a fair comparison across different schedulers, we augment all choices of $f(.)$ to $f^{K}(.)$ such that $\lfloor f^{K}(t_{k-1})L\rfloor<\lfloor f^{K}(t_{k})L\rfloor$ (i.e. at least one token is decoded at each step). Note that in the sampling process, the exact real value of $f^{K}(t_{k})$ does not matter as long as it does not change $\lfloor f^{K}(t_{k})L\rfloor$. Hence, we can parameterize the sampling process in an alternative manner using a sequence of integers $F^{K}_{k}=\lfloor f^{K}(t_{k})L\rfloor$, with $F^{K}_{0}=0$ and $F^{K}_{1}=L$. Formally, set $F^{K}$ by solving the following optimization objective
@@ B.4 Prefix-DLM
Table 11: Ablation Studies on the Choice of Vision Encoders. We compare the performance of models with different vision encoders after 1,500 steps of training.
@@ Appendix D Limitations
We also tried to extend the context length of these models via techniques such as rope rescaling, but achieved limited success. We experimented with extending Dream to 4096 context length and evaluate the model using the needle-in-a-haystack task at 4096 context length. We found that Dream-7B (72.4 Acc) performs worse than LLaDa-8B (91.6 Acc) and Llama-3-8B (95.4 Acc).
@@
Appendix E Reproducibility Statement
We will release the model checkpoints, training and inference code, as well as the training data to reproduce our results.
Appendix F Boarder Impacts and Safeguards
@@ Appendix H Concurrent Work
Concurrent to this work, MMaDa[77] a unified understanding and generation modeling using DM formulation. However, they do not leverage many of the novel techniques that we propose, leading to inferior performance and slow inference speed. We list a brief comparison in Table 12.

### v2

@@ A.1 Formulation of Discrete Diffusion Models (DMs)
In this section, we provide a detailed review of the formulation of Discrete Diffusion Models (DMs) briefly described in 2.2. Given a text sequence of $L$ tokens $X_{0}=[X_{0}^{1},X_{0}^{2},...X_{0}^{L}]$, the forward process $q(X_{t}|X_{s})$ gradually coverts it to a sequence of mask tokens $[M]$, denoted by $X_{1}=[X_{1}^{1},X_{1}^{2},...X_{1}^{L}]$ over the continuous time-interval $[0,1]$, with $1\geq t\geq s\geq 0$. Formally, this process is defined as
@@ A.2 Timestep shifting Schedule
Rounding in Discretization. In principle, we can pick any $f(.)$. However, given a particular choice of $L$ and $K$, if $\lfloor f(t_{k})L\rfloor$ and $\lfloor f(t_{k-1})L\rfloor$ yields the same integer, then no tokens are unmasked when we compute $p_{\theta}(X_{t_{k-1}}|X_{t_{k}})$. Hence, the actual number of function calls to the model $\theta$ may be less than, $K$ depending on the choice of $f(.)$. To make the sampling compute cost more predictable and allow for a fair comparison across different schedulers, we augment all choices of $f(.)$ to $f^{K}(.)$ such that $\lfloor f^{K}(t_{k-1})L\rfloor<\lfloor f^{K}(t_{k})L\rfloor$ (i.e. at least one token is decoded at each step). Note that in the sampling process, the exact real value of $f^{K}(t_{k})$ does not matter as long as it does not change $\lfloor f^{K}(t_{k})L\rfloor$. Hence, we can parameterize the sampling process in an alternative manner using a sequence of integers $F^{K}_{k}=\lfloor f^{K}(t_{k})L\rfloor$, with $F^{K}_{0}=0$ and $F^{K}_{1}=L$. Formally, we set $F^{K}$ by solving the following optimization objective
@@ B.4 Prefix-DLM
Table 11: Ablation Studies on the Choice of Vision Encoders. We compare the performance of models with different vision encoders after 1,500 steps of training. On Average, SigLip improves over CLIP by +1.86 and improves over MetaCLIP by +2.77.
@@ Appendix D Limitations
We also tried to extend the context length of these models via techniques such as rope rescaling, but achieved limited success. We experimented with extending Dream to 4096 context length and evaluate the model using the needle-in-a-haystack task [28] at 4096 context length. We found that Dream-7B (72.4 Acc) performs worse than LLaDa-8B (91.6 Acc) and Llama-3-8B (95.4 Acc).
@@
Appendix E Boarder Impacts and Safeguards
@@
Appendix F License Information for Data and Models
We report the licenses of the used artifacts in Table 12. We followed the intended use of all respective artifacts.
Table 12: Licenses and sources for datasets and models used. Datasets marked with * have no published license; the accompanying paper’s arXiv license is used as fallback. Training entries marked with or reflect compound or inferred licensing conditions.
Category Name License Platform
Vision Encoder SigLIP Apache 2.0 Hugging Face
Language Model LLaDA-8B MIT Hugging Face
Language Model Dream-7B Apache 2.0 Hugging Face
VLM LLaVA-NeXT (1.6) Apache 2.0 GitHub
VLM LLaVA-NeXT (1.6) LLaMA 2 License Hugging Face
VLM Open-LLaVA-NeXT Apache 2.0 Hugging Face
Train Dataset LLaVA-Pretrain CC3M + BSD-3-Clause Hugging Face
Train Dataset Open-LLaVA-NeXT-mix1M CC BY-NC 4.0 Hugging Face
Evaluation Dataset MME (MME-P) arXiv Non-exclusive* arXiv
Evaluation Dataset VQAv2 CC BY 4.0 Official Website
Evaluation Dataset MMBench Apache 2.0 GitHub
Evaluation Dataset MMMU Apache 2.0 Hugging Face
Evaluation Dataset MME (MME-C) arXiv Non-exclusive* arXiv
Evaluation Dataset MathVista CC BY-SA 4.0 Hugging Face
Evaluation Dataset MathVerse MIT Hugging Face
Evaluation Dataset MathVision MIT Hugging Face
Evaluation Dataset ScienceQA CC BY-NC-SA Official Website
Evaluation Dataset AI2D arXiv Non-exclusive* arXiv
Evaluation Dataset TextVQA CC BY 4.0 Official Website
Evaluation Dataset DocVQA arXiv Non-exclusive* arXiv
Evaluation Dataset ChartVQA GPL-3.0 GitHub
Evaluation Dataset InfoVQA arXiv Non-exclusive* arXiv
Note : Subject to the CC-3M dataset may be freely used for any purpose, although acknowledgement of Google LLC ("Google") as the data source would be appreciated. The dataset is provided "AS IS" without any warranty, express or implied. Google disclaims all liability for any damages, direct or indirect, resulting from the use of the dataset
Note : Our training data is based on Open-LLaVA-NeXT-mix1M, which combines ShareGPT4V data (CC BY-NC 4.0, research-only, with restrictions from LLaMA, Vicuna, and GPT-4 licenses) and AVG-LLaVA (Apache 2.0). All data was used strictly for academic research.
@@ Appendix H Concurrent Work
Concurrent to this work, MMaDa[78] (released after submission deadline) proposed a unified understanding and generation modeling using DM formulation. However, they do not leverage many of the novel techniques that we propose, leading to inferior performance and slow inference speed. We list a brief comparison in Table 13.